| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  177  
 | 
Sat Apr  9 04:01:35 2022 | 
Carsten |   | 2nd DAQ | Switched the CAEN TDC back on and changed 
the threshoolds for the ADC to: 
 
 | 
   | 
| 
  176  
 | 
Sat Apr  9 03:57:10 2022 | 
Carsten | Accelerator | New Cooler voltage after the crash of the cooler | After the restart of the cooler the Ohmlabs 
divider showed a voltage 16V lower than before. 
This could also be seen as a slight shift | 
   | 
| 
  175  
 | 
Sat Apr  9 03:46:31 2022 | 
Carsten | Accelerator | Slits in TE line | Changed GTEDSAH Rechts to 10mm | 
   | 
| 
  174  
 | 
Sat Apr  9 03:27:47 2022 | 
Patrick | DAQ | Laser scanning | DAQ1: lxg1927:/data.local2/BeamTimeTh22/
 
DAQ2: lxg0155:/data.local2/E142/
 
LMD file name DAQ1:229Thor
 | 
    | 
| 
  173  
 | 
Sat Apr  9 02:55:38 2022 | 
Carsten | Accelerator | Kühlerausfall und Gründe dazu | Wir hatten heute Abend eine Kühlerausfall, 
der uns mehrere Stunden beschäftigt hat, 
weil  gem. Anzeige in Device Control kein | 
   | 
| 
  172  
 | 
Fri Apr  8 23:56:07 2022 | 
Patrick | DAQ | DAQ improvement | As requested in ID:146 and ID:147, the laserscan 
tab was improved:
 
[LIST]
 | 
   | 
| 
  171  
 | 
Fri Apr  8 22:26:44 2022 | 
Patrick | DAQ | Spike in all detectors | In the previous shift, there was a spike in 
all detectors in the same scan step. Apparently, 
this was NOT a resonance since in the next | 
   | 
| 
  170  
 | 
Fri Apr  8 17:57:45 2022 | 
Konstantin | DAQ | Laser scanning | 
 
 
[TABLE border="1"]
 | 
    | 
| 
  169  
 | 
Fri Apr  8 17:44:25 2022 | 
Konstantin | Calibration | Final Timing | We measured the timing for the laser bunch 
synchronization. A summary is given below.
 
 | 
   | 
| 
  168  
 | 
Fri Apr  8 15:45:24 2022 | 
YL |   | Spectrum manipulation 5 | See attached | 
    | 
| 
  167  
 | 
Fri Apr  8 15:42:20 2022 | 
Yuri | Accelerator | Slits in the TE line | At around 15:00, The slits in the TE line 
were changed to :
 
TE4DS5HL -12 mm -> -20 mm
 | 
   | 
| 
  166  
 | 
Fri Apr  8 15:39:36 2022 | 
all | Runs | first data after timing set up | started taking data again
 
 
started at 15:40  
 | 
   | 
| 
  165  
 | 
Fri Apr  8 14:46:02 2022 | 
Danyal & Sebastian & Volker & Ken | Detectors | high dark countrate Cu PMT solved | The problem - "discovered" by Volker and Ken 
- of the high dark countrate of the Cu PMT 
has been solved. 
 | 
     | 
| 
  164  
 | 
Fri Apr  8 14:43:50 2022 | 
Wilfried | DAQ | laser timing - Q-switch start | below are my notes concerning laser timing 
/ Q-switch start | 
    | 
| 
  163  
 | 
Fri Apr  8 14:14:28 2022 | 
Kristian | DAQ | Delay of bunch RF | The bunch HF is now fed through a delay to 
shift the stop trigger relative to the PMT 
signal. The delay generator is placed directly | 
   | 
| 
  162  
 | 
Fri Apr  8 14:12:57 2022 | 
morning shift | DAQ | timing for h=1 | The timing for h=1 has been set up 
 
 
The time window is now twice as large as | 
    | 
| 
  161  
 | 
Fri Apr  8 12:04:13 2022 | 
Ragan | Runs | NTCAP started | NTCAP is started for the data taking.
 
 
Carrier Frequency: 242.02 M
 | 
    | 
| 
  160  
 | 
Fri Apr  8 11:16:57 2022 | 
Kristian | Laser | Wavemeter | We changed the fiber to the wavemeter. The 
third fiber worked best and we could operate 
the WM at low exposure times, i.e., below | 
     | 
| 
  159  
 | 
Fri Apr  8 11:02:04 2022 | 
DW, YL, RSS, SL |   | Change h=2 to h=1 | After changing to h=1
 
The frequency manipulation #5: f=242.021487MHz 
(124th harmonic) [no bunching]
 | 
   | 
| 
  158  
 | 
Fri Apr  8 10:44:55 2022 | 
Ragan, Yuri, Danyal, Sergey | Runs | NTCAP stopped | NTCAP is stopped. Sergey will now set an ion 
bunch instead of two in the ESR.
 
 | 
   |